OpenStudy (anonymous):

Solve the equation. -7(2x - 1) = 35 x =

hero (hero):

Hint: Divide both sides by -7 to get: 2x - 1 = -5 Now finish solving for x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Solve the equation. 43 = 7w - 15 - 5w w =

OpenStudy (anonymous):

43 = 2w - 15 => 2w = 43 - 15 = 28 => w = 14.